OpenStudy (anonymous):

Factor completely 49x*2 − 81.

pooja195 (pooja195):

Diffrence of squares \[\huge~\rm~\sqrt{49x^2}-\sqrt{81}\]

OpenStudy (danjs):

Try , since both coefficients are squares, and a minus sign.....( ( a + b) * ( a - b)

OpenStudy (anonymous):

@pooja195 that setup is much more easier.

pooja195 (pooja195):

What set up?

OpenStudy (anonymous):

When you square root the numbers instead of trying to factor them. Some people don't think of square rooting the numbers.

pooja195 (pooja195):

Right :) thanks ^-^ @DanJS 's work is helpful too since it belongs in that form :)

OpenStudy (anonymous):

is it (7-9)*2

OpenStudy (danjs):

I think that is what it looks like, if you have perfect squares a and b \[a^2*x^2 - b^2 = (a+b)*(a-b\]

OpenStudy (danjs):

sorry (ax - b)*(ax+b)

OpenStudy (danjs):

distribute backwards and check, it works out both ways

pooja195 (pooja195):

No rember the x's...u have it some what correct (7x+9) (7x-9) We need a positive and a negative statement